OpenStudy (anonymous):

Solve for d. -10 = -2 +1/2d

OpenStudy (texaschic101):

first add 2 to both sides

OpenStudy (texaschic101):

then divide both sides by 1/2

OpenStudy (anonymous):

-20

OpenStudy (texaschic101):

wrong answer Plainejane

OpenStudy (anonymous):

I believe the answer is x = -16.. But I am not sure.

OpenStudy (anonymous):

* d = -16

OpenStudy (anonymous):

Was I correct? @ysenia :)

OpenStudy (texaschic101):

-10 = -2 + 1/2d -10 + 2 = 1/2d -8 = 1/2d -8/(1/2) = d -8 * 2/1 = d -16 = d check... -10 = -2 + 1/2(-16) -10 = -2 - 8 -10 = -10 (correct) ANSWER : d = - 16

OpenStudy (texaschic101):

you are correct mercedes

OpenStudy (anonymous):

Thanks, Glad I could help. :)